OpenStudy (anonymous):

what is the lcd of 7,8 and 3

OpenStudy (anonymous):

168

OpenStudy (anonymous):

Since there are no common factors, just multiply the three numbers: 3*7*8

jimthompson5910 (jim_thompson5910):

7: 7*1 8: 2*2*2 3: 3*1 Circle all unique factors that appear the most. These factors are: 2*2*2, 3 and 7. Multiply them all out to get 2*2*2*3*7 = 8*21 = 168 So the LCM is 168